please help me solve.three times the sum of -4 and a number yields the difference between 20 and 5 times the number.
Answer by Theo(13342)

x is the number.
3 * (-4 + x) = 20 - 5x
simplify to get -12 + 3x = 20 - 5x
add 12 to both sides and add 5x to both sides to get:
3x + 5x = 20 + 12
combine like terms to get 8x = 32
divide both sides by 8 to get x = 4

the following article talks about translating key words into algebraic statements that should be helpful.

http://www.purplemath.com/modules/translat.htm

be sure to look at both pages, since the second page has examples of how to do the translations.

the order of the operations is important.

in this problem it didn't matter since both sides of the equation became 0.

that doesn't happen all the time, so make sure you translate the words into the equation in the right order.